OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point given for the correct solution. x 2 ---------- + ---- =0 x+3 5

OpenStudy (hoblos):

x -2 ----------- = -------- x+3 5 cross multiply 5x = -2x -6 7x = -6 x = -6/7